# A General Simulation-Based Optimisation Framework for Multipoint Constant-Stress Accelerated Life Tests

### v1

@@ 1 Introduction
The reliability of many modern components is such that it can be quite difficult to obtain failure information using standard life tests. In other words, the time-to-failure of reliable components is often so large it is not practically feasible to obtain lifetime data under normal operating conditions. Accelerated life testing (ALT) is a method of subjecting components to extreme stress to expediate, i.e. accelerate, their failure time. ALT can be used to discover failure modes that may occur in normal use, but more typically, the goal is to use the accelerated (high stress) data to predict the component lifetime under intended operating conditions (design stress) (Nelson,, 2005). This involves fitting a statistical model to the accelerated data from which extrapolations are made to estimate the lifetime at the design stress.
@@
The proposed simulation framework allows a variety of test plans to be evaluated and compared offline, thus offering a more time and cost efficient alternative to physical (destructive) component testing. By generating random lifetime data, we can compare the true (known) parameters against values estimated from statistical models for a given test plan configuration. Within this simulation framework we can modify the number of design points (stresses) and their levels, alter the test unit allocation, and explore various life-stress relationships to assess a range of test plans in terms of their design-stress lifetime prediction performance. We optimise accelerated test plans by employing the differential evolution (DE) optimisation algorithm, where the inputs to the objective function are the test plan parameters, and the optimisation criterion is the RMSE (root mean squared error) of out-of-sample lifetime predictions.
@@
The task of obtaining an optimal accelerated test plan has a wide literature, with several published works focussing on specific test scenarios, i.e., test plans for a given lifetime distribution, stress type and number of stresses. For example, a two-stress, constant-stress test plan is considered by Barton, (1991). Meeker and Nelson, (1975) and Nelson and Meeker, (1978) also consider two-stress constant-stress plans in relation to the Weibull and extreme value distributions, whilst Yang, (1994) discusses a constant-stress test plan with up to four stress levels. Step-stress test plans are considered by Haghighi, (2014), and Khamis and Higgins, (1996) discuss a three-step step-stress test for both linear and quadratic models. More recently, Ayasse and Seo, (2022) demonstrate the optimisation of an accelerated life test with two stress levels using simulated data and response surface methodology.
@@
In contrast to much of the existing literature, we do not focus on a specific ALT scenario, but instead propose a general and flexible framework that is both applicable across a range of ALT optimisation problems, and compatible with various common physics-based/empirical life-stress relationships. For simpler experiments, with a limited number of design points, it is possible to mathematically optimise the test plan, and we find that our simulation results are consistent with existing analytical results in such scenarios. However, for complex multipoint designs, theoretical/analytical optimisation is more challenging, and to the best of our knowledge, such analyses are lacking in the literature. Where theoretical optimisation could prove intractable in such multipoint design scenarios, our proposed optimisation procedure can easily accommodate design schemes with an arbitrarily large number of design points. Furthermore, our flexible statistical modelling framework permits any statistical distribution (with any model functional form). Therefore, complex life-stress relationships can also be incorporated within complex design schemes. This procedure offers a practical alternative to theoretical analysis and serves as an effective tool for optimising complex multipoint accelerated life test plans. Our aim is to provide actionable insight to reliability practitioners working on real-life ALT problems.
@@
The remainder of this article is organised as follows. In Section 2 we discuss test plan design considerations. Details of the modelling and simulation framework are provided in Section 3, where our optimisation procedure is applied in some commonly-used settings (linear, quadratic and power-law stress) in Section 4. In Section 5, we utilise a real-world ALT case study to demonstrate the application of our optimisation procedure. Finally, the main findings and results are discussed in Section 6.
@@ 2 Test Plan Considerations
After conducting an accelerated life test, a statistical model is fitted to the (high stress) data from which extrapolations are made to predict the lifetime at the design stress (intended operating stress). The purpose of the test plan is to facilitate the model fitting process by extracting as much information as possible from the experiment. In designing the test plan, the main considerations include: (i) accelerating variable(s), (ii) stress loading, (iii) test duration, (iv) lifetime distribution, (v) life-stress relationship and (vi) optimality criterion. The extensive reviews from Chen et al., (2018), Limon et al., (2017) and Nelson, (2005) draw the conclusion that the majority of test plans utilise a single (constant stress) accelerating variable and Type I data where components are tested until the earlier of failure or a specified testing duration, and an assumed Weibull lifetime distribution.
We aim to provide a practical optimisation framework that supplements physical ALT experiments. Therefore, in this work, we focus on the experimental framework most likely to be adopted in practice: Type I data, single-variable, constant-stress test plans. In relation to the optimality criterion, most works aim to minimise the asymptotic variance of the maximum likelihood estimate (MLE) of an estimated quantile for a specified low stress (Nelson,, 2005; Chen et al.,, 2018). In this work, we consider both estimation variability and bias using the RMSE of the estimated quantile, where we numerically approximate the RMSE through simulation. In terms of specifying a lifetime distribution, we align with the ALT literature, and focus on the Weibull distribution. However, as outlined in the following section, our modelling framework is sufficiently general to accommodate models with any underlying lifetime distribution, and with any lifetime-stress relationship.
@@ 3 Modelling Framework
Estimating the design-stress lifetime requires a statistical model that faithfully describes the underlying lifetime-stress relationship. Once such a model is obtained, one can extrapolate the model to predict the lifetime at the design stress. The optimal test plan is that which maximises the lifetime information, ultimately improving the prediction performance of the fitted model. In this section we detail the statistical modelling framework and outline our simulation-based approach to obtaining optimal test plans.
@@ 3.1 Statistical Model
Letting $s_{d}$ be the design stress, in this paper, we focus on the median design-stress lifetime, $q_{0.5}(s_{d})$.
In ALT, a life-stress relationship is often specified from empirical or physics-based models. For example, when temperature is used as the stress condition the Arrhenius relationship can be used (Ayasse and Seo,, 2022; Limon et al.,, 2017). When the accelerating variable is voltage, the inverse power law model is often used (Barton,, 1991; Meeker and Nelson,, 1975), with the generalised Eyring model being a popular choice to capture the combined effects of thermal and non-thermal stresses (Limon et al.,, 2017). However, in this work, we adopt a very general modelling approach, utilising an accelerated failure time (AFT) model of the form
@@
The comprehensive bibliographies of Chen et al., (2018) and Limon et al., (2017) show that in terms of specifying a life-stress relationship, the majority of references use the exponential, log-normal and Weibull distributions. Although the modelling framework outlined above can easily accommodate any probability distribution (through choice of the distribution of $\epsilon$), in this paper, we focus on the most widely used distribution, the Weibull. Given $T\sim\text{Weibull}$, the median design-stress lifetime is given as
@@ 3.2 Optimisation Framework
Denoting $N$ as the number of stresses (design points), $\bm{S}=\{S_{1},S_{2},\cdots,S_{N}\}$ as the stress levels, $\bm{P}=\{p_{1},p_{2},\cdots,p_{N}\}$ as the proportion of test units allocated to a given stress level ($p_{1}+p_{2}+\cdots+p_{N}=1$), and $Q=q_{0.5}(s_{d})$, the objective function of interest (which we aim to minimise) is given by
$\text{RMSE}(N,\bm{S},\bm{P})=\left(E[\hat{Q}-Q]^{2}\right)^{1/2},$
where $\hat{Q}$ is an estimate of $Q$. This RMSE objective function depends on the configuration of the testing experiment, i.e., the test plan parameters: N, $\bm{S}$, $\bm{P}$, as a direct consequence of $\hat{Q}$ depending on the test plan, whereas $Q$ is a fixed constant. Thus, in general, particularly for complex design scenarios, it may prove infeasible to compute the RMSE analytically. Therefore, in order to provide a sufficiently general optimisation framework that accommodates complex designs, we propose approximating the RMSE by way of Monte Carlo simulation, i.e.,
$\text{RMSE}(N,\mathbf{S},\mathbf{P})\approx\left(\dfrac{1}{n_{\text{sim}}}\displaystyle\sum_{j=1}^{n_{\text{sim}}}[\hat{Q_{j}}-Q]^{2}\right)^{1/2},$ (2)
where $n_{\text{sim}}$ is the number of simulation replicates; in this paper, we use $n_{\text{sim}}=1000.$
Given that the RMSE objective depends on the inputs ($N,\mathbf{S},\mathbf{P}$) in a complex non-linear (and possibly non-differentiable) manner, we use the differential evolution (DE) optimisation algorithm. The DE algorithm is a type of genetic algorithm that adopts a biological, mutation-based approach to evolve the objective function over successive iterations (generations). (See Storn and Price, (1997) for further details.) Within our optimisation framework, we specifically utilise the DE algorithm from the R package DEoptim, developed by Mullen et al., (2011). The DE algorithm has been shown to perform well across a wide range of situations where the objective function is noisy/discontinuous and/or difficult to differentiate (Mullen et al.,, 2011). Additionally, Price et al., (2005) demonstrate the versatility (and robustness) of the algorithm in the context of various practical problems. Thus, the DE algorithm, and its R implementation, DEoptim, is an appropriate choice for our general modelling and test-plan setup with stochastic RMSE objective function.
@@ 3.3 Parameter Pre-Estimation
As mentioned in the previous section, to produce an optimal test plan in practice, an initial estimate of model parameters is needed in order to generate accelerated data that is consistent with the underlying component lifetime distribution. This problem of requiring knowledge of the true model parameters is inherent in all accelerated life test plan optimisation (Nelson and Meeker,, 1978; Nelson,, 2005). Chen et al., (2018) also acknowledges that initial model parameter estimates are required to design an optimal ALT test plan. One solution is to use a Bayesian analysis approach where unknown values are replaced with prior distributions. Examples of Bayesian approaches in ALT planning include Chaloner and Larntz, (1992), Naijun and Teng, (2017) and Zhang and Meeker, (2006).
Another common approach is to conduct a pilot study/preliminary test and/or utilise lifetime information from a similar component. For example, in their discussion of optimising a two-stress temperature-stress test plan, Ayasse and Seo, (2022) use historical data from a similar product to obtain log-normal model parameter estimates. Barton, (1991) similarly provides a temperature-based example (using the Arrhenius model) that utilises preliminary test data. In a discussion on assumed parameter values, Nelson and Meeker, (1978) state that, in practice, parameters must be estimated using experience, similar data, or a preliminary test. Yang, (1994) too discusses the issue of pre-estimating unknown parameters and suggests the same solutions as Nelson and Meeker, (1978), but also mentions the Bayes approach. In this work, we adopt the practical and commonly-used approach of obtaining initial parameter estimates from a preliminary testing experiment.
@@ 3.4 Preliminary Test Data
Our approach is to fit a suitable reliability model to preliminary test data, and use the fitted model coefficients as the parameters in the optimisation. We consider a preliminary test where $n$ components, each with time-to-failure, $T$, are subjected to a single (constant) accelerating stress, $S\in(0,1)$, such that $T|(S=0)=\infty$, and $T|(S=1)=0$. In practice, a preliminary experiment may have physical constraints such as testing equipment capacity, component availability etc., and as such will be application specific. For demonstration purposes, in the next section, we consider $n=180$ components tested at the following stresses: $S=\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$, and specify a design-stress of $s_{d}=0.05$. In standard constant-stress tests, it is common to use equally spaced stresses, with an equal number of test units at each stress (Tang et al.,, 2002), so for this initial test, we consider a balanced design, i.e., $n_{1}=n_{2}\cdots=n_{9}=20$. In this paper, we focus on preliminary test data which follows a Weibull distribution, and consider some common statistical functional forms, specifically: linear, quadratic and power law relationships. In the following section, we give a brief overview of the optimisation schemes considered, and discuss the results of each optimisation in detail.
@@ 4 Optimisation Results
(i) For each fixed $N\in\{2,3,4,5,6\}$, determine the optimal $N$-point design. $N$ is specified and the objective function inputs are the stress levels, $\mathbf{S}$, and test units at each stress, given by $n\times\mathbf{P}$. (Of course, larger values of $N$ can be used within our framework.)
@@
Note: Given the large number of test plans analysed in this work (each with a large number of simulation replicates), in the interest of reducing computational cost, we fix the number of optimisation iterations (DEoptim generations) to be 50. We have found, that for this application, 50 iterations offers a good compromise between computational efficiency and stability of the optimal plan. However, this parameter is easily adjusted within the optimisation framework, hence one can increase (or decrease) the number of iterations to suit their specific application/requirements. It could be set smaller if fast initial plans are needed (e.g., in a dynamic visualisation tool) or larger for a final optimisation.

### v2

@@ 1 Introduction
The reliability of many modern components is such that it can be quite difficult to obtain failure information using standard life tests. In other words, the time-to-failure of reliable components is often so large it is not practically feasible to obtain lifetime data under normal operating conditions. Accelerated life testing (ALT) is a method of subjecting components to extreme stress to expediate, i.e. accelerate, their failure time. ALT can be used to discover failure modes that may occur in normal use, but more typically, the goal is to use the accelerated (high stress) data to predict the component lifetime under intended operating conditions (design stress) (Nelson:2005). This involves fitting a statistical model to the accelerated data from which extrapolations are made to estimate the lifetime at the design stress.
@@
The proposed simulation framework allows a variety of test plans to be evaluated and compared offline, thus offering a more time and cost efficient alternative to physical (destructive) component testing. By generating random lifetime data, we can compare the true (known) parameters against values estimated from statistical models for a given test plan configuration. Within this simulation framework we can modify the number of design points (stresses) and their levels, alter the test unit allocation, and explore various life-stress relationships to assess a range of test plans in terms of their design-stress lifetime prediction performance. Using the statistical computing language R, we optimise accelerated test plans by employing the differential evolution (DE) optimisation algorithm (and R’s DE package DEoptim), where the inputs to the objective function are the test plan parameters. For the purpose of this paper, we use the RMSE (root mean squared error) of out-of-sample lifetime predictions as our optimisation criterion. However, the overall framework, based on data simulation and DE optimisation, is sufficiently general such that any other optimality criterion and/or experimental constraints could be included.
@@
The task of obtaining an optimal accelerated test plan has a wide literature, with several published works focussing on specific test scenarios, i.e., test plans for a given lifetime distribution, stress type and number of stresses. For example, a two-stress, constant-stress test plan is considered by Barton:1991. Meeker:1975 and Nelson:1978 also consider two-stress constant-stress plans in relation to the Weibull and extreme value distributions, whilst Yang:1994 discusses a constant-stress test plan with up to four stress levels. Step-stress test plans are considered by Haghighi:2014, and Khamis:1996 discuss a three-step step-stress test for both linear and quadratic models. More recently, Ayasse:2022 demonstrate the optimisation of an accelerated life test with two stress levels using simulated data and response surface methodology.
@@
In contrast to much of the existing literature, we do not focus on a specific ALT scenario, but instead propose a general and flexible framework that is both applicable across a range of ALT optimisation problems, and compatible with various common physics-based/empirical life-stress relationships. For simpler experiments, with a limited number of design points, it is possible to mathematically optimise the test plan, and we find that our simulation results are consistent with existing analytical results in such scenarios. However, for complex multipoint designs, theoretical/analytical optimisation is more challenging, and to the best of our knowledge, such analyses are lacking in the literature. Where theoretical optimisation could prove intractable in such multipoint design scenarios, our proposed optimisation procedure can easily accommodate design schemes with an arbitrarily large number of design points. Furthermore, our flexible statistical modelling framework permits any statistical distribution (with any model functional form). Therefore, complex life-stress relationships can also be incorporated within complex design schemes. Our aim is to provide actionable insight to reliability practitioners working on real-life ALT problems. As such, the optimisation framework offers a practical alternative to theoretical analysis and serves as an effective tool for optimising complex multipoint accelerated life test plans. We also make our tool available in the form of open-source R code available at: https://github.com/o-mcgrath/ALT_Optimisation
@@
The remainder of this article is organised as follows. In Section 2 we discuss test plan design considerations. Details of the modelling and simulation framework are provided in Section 3, and our optimisation procedure is applied in some commonly-used settings (linear, quadratic and power law stress) in Section 4. In Section 5, we utilise a real-world ALT case study to demonstrate the application of our optimisation procedure. Finally, the main findings and results are discussed in Section 6.
@@ 2 Test Plan Considerations
After conducting an accelerated life test, a statistical model is fitted to the (high stress) data from which extrapolations are made to predict the lifetime at the design stress (intended operating stress). The purpose of the test plan is to facilitate the model fitting process by extracting as much information as possible from the experiment. In designing the test plan, the main considerations include: (i) accelerating variable(s), (ii) stress loading, (iii) test duration, (iv) lifetime distribution, (v) life-stress relationship and (vi) optimality criterion. The extensive reviews from Chen:2018, Limon:2017 and Nelson:2005 draw the conclusion that the majority of test plans utilise a single (constant stress) accelerating variable and Type I data where components are tested until the earlier of failure or a specified testing duration, and an assumed Weibull lifetime distribution.
We aim to provide a pragmatic optimisation framework that supplements physical ALT experiments. Therefore, in this work, we focus on the experimental framework most likely to be adopted in practice: Type I data, single-variable, constant-stress test plans. In relation to the optimality criterion, most works aim to minimise the asymptotic variance of the maximum likelihood estimate (MLE) of an estimated quantile for a specified low stress (Nelson:2005; Chen:2018). In this work, we consider both estimation variability and bias using the RMSE of the estimated quantile, where we numerically approximate the RMSE through simulation. In terms of specifying a lifetime distribution, we align with the ALT literature, and focus on the Weibull distribution. However, as outlined in the following section, our modelling framework is sufficiently general to accommodate models with any underlying lifetime distribution, and with any lifetime-stress relationship.
@@ 3 Modelling Framework
Estimating the design-stress lifetime requires a statistical model that faithfully describes the underlying lifetime-stress relationship. Once such a model is obtained, one can extrapolate the model to predict the lifetime at the design stress. The optimal test plan is that which maximises the lifetime information, ultimately improving the prediction performance of the fitted model. In this section we detail the statistical modelling framework and outline our simulation-based approach to obtaining optimal accelerated test plans.
@@ 3.1 Statistical Model
Letting $s_{d}$ be the design stress, in this paper, we focus on obtaining the median design-stress lifetime, $q_{0.5}(s_{d})$. However, the proposed framework permits any quantile (or indeed any quantity) of interest to be considered.
In ALT, a life-stress relationship is often specified from empirical or physics-based models. For example, when temperature is used as the stress condition the Arrhenius relationship can be used (Ayasse:2022; Limon:2017). When the accelerating variable is voltage, the inverse power law model is often used (Barton:1991; Meeker:1975), with the generalised Eyring model being a popular choice to capture the combined effects of thermal and non-thermal stresses (Limon:2017). However, in this work, we adopt a very general modelling approach, utilising an accelerated failure time (AFT) model of the form
@@
The comprehensive bibliographies of Chen:2018 and Limon:2017 show that in terms of specifying a life-stress relationship, the majority of references use the exponential, log-normal and Weibull distributions. Although the modelling framework outlined above can easily accommodate any probability distribution (through choice of the distribution of $\epsilon$), in this paper, we focus on the most widely used distribution, the Weibull. Given $T\sim\text{Weibull}$, the median design-stress lifetime is given as
@@ 3.2 Optimisation Framework
Denoting $N$ as the number of stresses (design points), $\bm{S}=\{S_{1},S_{2},\cdots,S_{N}\}$ as the stress levels, $D$ as the minimum allowable distance between consecutive stresses, $\bm{P}=\{p_{1},p_{2},\cdots,p_{N}\}$ as the proportion of test units allocated to a given stress level ($p_{1}+p_{2}+\cdots+p_{N}=1$), $t_{e}\in(0,\infty)$ as the experiment duration and $Q=q_{0.5}(s_{d})$ as the median lifetime at the design stress, the objective function of interest (which we aim to minimise) is given by
$\text{RMSE}(N,\bm{S},D,\bm{P},t_{e})=\left(E[\hat{Q}-Q]^{2}\right)^{1/2},$
where $\hat{Q}$ is an estimate of $Q$. This RMSE objective function depends on the configuration of the testing experiment, i.e., the test plan parameters, $N$, $D$, $\bm{S}$, $\bm{P}$ and $t_{e}$, as a direct consequence of $\hat{Q}$ depending on the test plan, whereas $Q$ is a fixed constant. In general, particularly for complex design scenarios, it may prove infeasible to compute the RMSE analytically. Therefore, in order to provide a sufficiently general optimisation framework that accommodates complex designs, we propose approximating the RMSE by way of Monte Carlo simulation, i.e.,
$\text{RMSE}(N,\bm{S},D,\bm{P},t_{e})\approx\left(\dfrac{1}{n_{\text{sim}}}\displaystyle\sum_{j=1}^{n_{\text{sim}}}[\hat{Q_{j}}-Q]^{2}\right)^{1/2},$ (2)
where $n_{\text{sim}}$ is the number of simulation replicates. Of course, the larger the value of $n_{\text{sim}}$, the lower the Monte Carlo error, i.e., the closer (2) will be to the analytic (but intractable) RMSE. In this paper, we use $n_{\text{sim}}=1000$, but the results are very similar for $n_{\text{sim}}=500$ (see Appendix B.3).
Given that the RMSE objective depends on the inputs ($N,\bm{S},D,\bm{P},t_{e}$) in a complex non-linear (and possibly non-differentiable) manner, we use the differential evolution (DE) optimisation algorithm. The DE algorithm is a type of genetic algorithm that adopts a biological, mutation-based approach to evolve the objective function over successive iterations (generations). (See Storn:1997 for further details.) Within our optimisation framework, we specifically utilise the DE algorithm from the R package DEoptim, developed by Mullen:2011. The DE algorithm has been shown to perform well across a wide range of situations where the objective function is noisy/discontinuous and/or difficult to differentiate (Mullen:2011). Additionally, Price:2005 demonstrate the versatility (and robustness) of the algorithm in the context of various practical problems. Thus, the DE algorithm, and its R implementation, DEoptim, is an appropriate choice for our general modelling and test-plan setup with stochastically-approximated RMSE objective function. It is clear that our proposal is quite flexible: (i) an objective function of interest is computed, which could be RMSE (per this paper) or some other function; (ii) the computation is based on simulated data, wherein any realistic experimental constraints can be included, e.g., the experimental duration, the minimum distance between stresses, the stresses having to fall in a discrete set, or the number of test units having to be multiples of some number; and (iii) given the objective function and experimental constraints, the design is optimised using the general purpose DE algorithm.
@@ 3.3 Parameter Pre-Estimation
As mentioned in the previous section, to produce an optimal test plan in practice, an initial estimate of model parameters is needed in order to generate accelerated data that is consistent with the underlying component lifetime distribution. This problem of requiring knowledge of the true model parameters is inherent in all accelerated life test plan optimisation (Nelson:1978; Nelson:2005). Chen:2018 also acknowledge that initial model parameter estimates are required to design an optimal ALT test plan. One solution is to use a Bayesian analysis approach where unknown values are replaced with prior distributions. Examples of Bayesian approaches in ALT planning include Chaloner:1992, Naijun:2017 and Zhang:2006.
Another common approach is to conduct a pilot study/preliminary test and/or utilise lifetime information from a similar component. For example, in their discussion of optimising a two-stress temperature-stress test plan, Ayasse:2022 use historical data from a similar product to obtain log-normal model parameter estimates. Barton:1991 similarly provides a temperature-based example (using the Arrhenius model) that utilises preliminary test data. In a discussion on assumed parameter values, Nelson:1978 state that, in practice, parameters must be estimated using experience, similar data, or a preliminary test. Yang:1994 too discusses the issue of pre-estimating unknown parameters and suggests the same solutions as Nelson:1978, but also mentions the Bayes approach. In this work, we adopt the practical and commonly-used approach of obtaining initial parameter estimates from a preliminary testing experiment.
@@ 3.4 Preliminary Test Data
Our approach is to fit a suitable reliability model to preliminary test data, and use the fitted model coefficients as the parameters in the optimisation. We consider a preliminary test where $n$ components, each with time-to-failure, $T$, are subjected to a single (constant) accelerating stress, $S\in(0,1)$, such that $T|(S=0)=\infty$, and $T|(S=1)=0$. In practice, a preliminary experiment may have physical constraints such as testing equipment capacity, component availability etc., and as such, will be application specific. For demonstration purposes, in the next section, we consider $n=180$ components tested at the following stresses: $\bm{S}=\{0.1,0.2,\cdots,0.8,0.9\}$, and specify a design stress of $s_{d}=0.05$. In standard constant-stress tests, it is common to use equally spaced stresses, with equal test unit allocation (Tang:2002), so for this initial test, we consider a balanced design, i.e., $n_{1}=n_{2}\cdots=n_{9}=20$. In this paper, we focus on preliminary test data which follows a Weibull distribution, and consider some common statistical functional forms, specifically: linear, quadratic and power law relationships. In the following section, we give a brief overview of the optimisation schemes considered, and discuss the results of each optimisation in detail.
@@ 4 Optimisation Results
(i) For each fixed $N\in\{2,3,4,5,6\}$, determine the optimal $N$-point design. $N$ is specified and the objective function inputs are the stress levels, $\bm{S}$, the minimum distance between consecutive stresses, $D$, the test units at each stress, given by $n\times\bm{P}$ and the experiment duration, $t_{e}$. (Of course, larger values of $N$ can be used within our framework.)
@@
Note: In our analysis, the number of optimisation iterations (DEoptim generations) is fixed at 200 (DEoptim default), however, this parameter is easily adjusted within the optimisation framework. It is largely a matter of computational resources and time-sensitivity of results: it could be set smaller if fast initial plans are needed (e.g., in some initial exploratory work) or larger for a final optimisation.